If a person walk at 14km/hr instead of 10km/hr,he would have walked 20km more .the actual distance travelled by him is
Found 2 solutions by jorel555, ewatrrr:
Answer by jorel555(1290)

Let h be the number of hours walked. Then:
14h=10h+20
4h=20
h=5
10 k/h x 5 hours=50 kilometers. ☺☺☺☺

Answer by ewatrrr(24785)

10kmph*t + 20km = 18kmph*t
20 = 8t
20/8 = t = 2.5hr
10kmph(2.5hr) = 25km, distance walked at 10kmph
Checking
25km + 20km = 45km = 18(5/2)km